Calculate the molar mass of the following substances.
       
       (e) Nitric acid, HNO3

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

Nitric acid, HNO3= 1 × 1 + 1 × 14 + 3 × 16
= 1 + 14 + 48 = 63 g


Thanks